Feder Observatory

A Blog for Minnesota State University Moorhead's Paul P. Feder Observatory

Archive for the ‘Repair or Maintenance Report’


Observing Log, 4/1/2014 0

Posted on April 02, 2014 by Matt

I went observing this evening; conditions were excellent throughout the night (consistent seeing, though the focus needed to be adjusted around 3:45AM because the temperature had gone down).

 

Air temperature started around 23F, focus set to 2795. Focus was changed to 2785 around 3:45AM because it had cooled off.

Targets:

  • M82 (enough images to get SNR on the supernova, but not enough to make pretty pictures)
  • UZ Vir, discussed in this article by Sodor, et al., because it had a maximum early-ish in the evening (12:40PM) and I’d like to try and reproduce their results (by which I really mean make my curve line up with theirs).
  • EY UMa, because I’m here, it is up, and there was a maximum forecast around 5AM.
  • The standard field SA104

Troubles:

  • The filter wheel briefly stopped working because the set screw was too tight. I think that is because it has warmed up considerably since the last time it was adjusted and the wheel could no longer turn.
  • The dome wasn’t lined up with the scope perfectly. Fixed by manually steering the dome a bit and setting its azimuth to the telescope’s azimuth.
  • Batteries in the reticle are dead.
  • I forgot to connect the scope to MaxImDL, so header patching will be more fun than usual.

Other than that it was smooth sailing!

Telescope computer maintenance and experimentation 0

Posted on March 06, 2014 by Matt

I came to the observatory today to do a few things:

  • Update software
  • Try running ACP (astronomer’s control panel), a piece of software for automating some aspects of observing.
  • Peek at the wiring of the computer and telescope to better understand how things fit together.
  • Try to find the State Property ID tag on the computer.

Software updated:

  • AVG Free anti virus updated to 2014 version.
  • Updated AVG virus definition files.

There is a new version of MaxImDL (v5.24, we have v5.23) but the changes seem pretty minor and we’d have to pay for the upgrade.

After creating a system restore point in case something goes wrong, I installed ACP v7.1, FocusMax 3.8.0.8 and the GSC 1.1.

On the upside, it doesn’t break anything that currently works. On the downside, it doesn’t give us any new capabilities until the ASCOM/ACL driver issues are fixed. It must be possible to talk to the telescope via ACL because TheSky v5 does it successfully.

 

Tried TheSky-controlled telescope ASCOM driver, no luck. Apparently I would need to set up Sky v5 for scripting, didn’t get a chance to check how to do that.

And I have no idea where the property tag is…

Observing Log 11/12/2013 0

Posted on November 13, 2013 by Matt

It was an…interesting…evening.

It started out well enough with taking calibration images (floodlight flats with diffuser) in B, V, and R along with darks and biases.

The only hangup was that a windows process called svchost kept taking up 100% of the CPU time. It launched every time Windows tried to update. Killing the process fixed the problem.

Opening the dome and initial pointing went fine too. Then, like the night before, MaxIm DL hung. I ended up putting the telescope back at zenith and rebooting the system.

Temperature was 27F and focus set to 2800.

After that was another brief, trouble-free interlude of data taking–this time BVR images of the standard field SA95.

Then I tried slewing to EY UMa; the telescope decided the shortest route would be to rotate the telescope towards the ground…so the scope stopped moving when it hit its limits. I tried a few times to push the telescope back up (which was fine, it moved) and re-slew. Every time it tried to slew through the ground. Eventually I pushed it up to (roughly) zenith and tried slewing from there but by that point the telescope was completely confused as to what its position was. So, back up to zenith by hand, shut down TCS and restart it (fortunately the dome was fine, so before shutting down I could send it home).

After the restart slewing was fine, though the initial pointing was a little challenging because the finder was no longer well-aligned with the telescope. That has been fixed (and the reticle is working).

The rest of the night was pretty uneventful–observed EY UMa in BVR through to dusk.

As the sun rose it became apparent that it had gotten partly cloudy overnight; that was also apparent from the change in flux of one of the comparison stars used for EY UMa.

PS If you notice a lemon-y fresh sent in the control room that is because I cleaned the counters.

Observing log 11/11/2013 0

Posted on November 12, 2013 by Matt

This is the first frigid night of the year: forecast low is 5F. Focus was set to 2775 with FWHM around 6 pixels (though it varied quite a bit).

First, some bad news:

+ the finder scope eyepiece barely stays in and the threads seem stripped. I might have done that last time I was out–bumped it and it fell out. For now it is still in the telescope.
+ The reticle batteries are dead and there aren’t replacements out here. I’ll fix that on Nov 12. For some reason taking the batteries out and putting them back in made the reticle work again. Will still bring out more replacement batteries.

Better news:

+ You can position the telescope without any of that. Detailed instructions will be written up, but basically: take a deep image (say 60 sec), use astrometry.net to do an astrometric fit (can now be done from within AstroImageJ), then use the center of the field it finds as the telescope position.

MaxIm DL hung around 11:30PM CST…no idea why. Restarting MaxIm DL solved the problem.

Targets tonight:

+ UV20598
+ EY UMa
+ standard field SA95

All images were taken in the R band.

And, a bonus…video of a twinkling star from focusing: https://www.youtube.com/watch?v=Fj30-tzKpqg

Problems with times stamp on 9/11, 9/12, 9/13 0

Posted on September 16, 2013 by Matt

I realized tonight that there was a problem with the observatory computer; the program that keeps the computer’s internal clock synchronized with Internet time servers wasn’t working.

It turns out the problem had nothing to do with that program. The IP addresses of campus DNS servers had changed so DNS lookup was failing.

When DNS was reconfigured the synchronization program reset the computer’s internal clock, adjusting it by -46.089 seconds. In other words, 46.089 seconds was subtracted from the computer’s internal clock.

Prior to tonight (9/16/2013) the last synchronization with the time server was on 8/20/2013 at about 5:45AM.

The bottom line is that the timestamp on images taken on the dates below are likely all off by some amount between 0 and -46 sec, but there is no way to know (at least at the moment) how big that offset is. On any given night the time difference between images should be very accurate. It is only the “absolute” time that is incorrect.

Updated MaxImDL to v5.23 0

Posted on August 13, 2013 by Matt

This release had some bug fixes but nothing major. Haven’t noticed any difference in how it operates so far.

Flat box maintenance 0

Posted on August 12, 2013 by Matt

I removed the aluminum foil cover from the round flat box and replaced it with aluminum tape. This shouldn’t affect taking flats in any way, but should be sturdier and longer-lasting than the foil solution.

Flat box with new light shielding

Flat box with new light shielding

Observing log, Apr 28, 2013 0

Posted on April 29, 2013 by Matt

WARNING: DO NOT USE DATA FROM THIS NIGHT, or use with extreme caution. Turns out the dome was causing much more extensive problems than I originally thought–the EY UMa data from this night had “jumps” in its light curve, each of which happened right after the telescope was slewed from EY UMa to a standard field and back.

Planning to observe EY UMa, some standard fields, and MY Aql tonight.

Conditions are very good, with seeing around 6 pixels FWHM.

Ran into a problem with the dome that affected observations from about 11:15PM-12:35AM CDT. Alignment of the slit with the telescope was off, so noise increased a lot.

Fixed by:

  • Turning “Auto Dome” to off in TCS/Telescope…/Misc…
  • manually steering the dome to match telescope position
  • Initializing dome position to current telescope position under TCS/Telescope…/Initialization…
  • Turned “Auto Dome” back on

No other problems during the night, and conditions were good for the entire night.

Observing Log, Apr 13, 2013 0

Posted on April 14, 2013 by Matt

Juan Cabanela and I came out to the observatory tonight. The filter wheel was reinstalled, and works properly. I was able to reproduce the ERROR=1 problem using the old CAT5 cable that connected the filter to the telescope pier in the dome. Replacing that cable fixed the problem.

 

The alignment of the camera with the sky is a little bit different now than it was before–North is now nearly perfectly in the top of the image, east is to the right.

I observed a few different variable stars, and the Landolt standard field SA110. Conditions were surprisingly good until about midnight, when clouds started to roll in. I wrapped up observing and closed the dome for shooting calibration images at 12:30AM, anticipating that snow will start within an hour or two.

 

GREAT to be out observing again!

New light box, June 21, 2012 0

Posted on June 21, 2012 by Matt

New light box has been mounted and tested under ambient white lights in dome.

The box is smaller and lighter weight than previous box.

Mounting:

1. Keep the black curtain snapped onto the box, which is a blue plastic feed bucket.  There is an opening in the curtain where the two ends meet, but are not fastened together.  Place this opening near the weight bar on the telescope so the curtain fits around the bar.

2.  The bucket simply mounts on top of the telescope, which must be positioned near zenith.

3. After placing the bucket on the rim of the telescope tube, you may have to pull the curtain edges out from under the bucket.  The curtain should hang freely all around the bucket, making a seal around the edges.

4. Please store light box in as clean a place as possible.  Do not place items inside the bucket or on any of its translucent surfaces.  For now, the bucket is under the covers in the dome until I find a plastic bag to store it in.

 

Data thus far, taken solely with ambient white light on in dome (not floodlight, just ordinary dome light):

Through R filter, a 180 sec exposure gives mean intensity 39,000

“” V filter, intensity with 360 secs gives intensity ~ 20,000

“” I filter intensity with 60 sec exp gives ~43,000



↑ Top